Rana is 25 years younger than his father. In 10 years time from now, his father will be twice Rana's age. Work out Rana's age. Suppose Rana's age now is x
Answer by AnlytcPhil(1806)

>>Suppose Rana's age now is x<<
>>Rana is 25 years younger than his father.<<
Therefore his father age is now is x + 25.
>>In 10 years time from now,...<<
Rana's age 10 years from now = x + 10
Father's age 10 years from now= Father's age now + 10 = (x + 25) + 10 = x+25+10 = x+35
>>his father will be twice Rana's age.<<
x + 35 = 2(x + 10)
x + 35 = 2x + 20
15 = x
Rana is 15 and his father is 15+25 or 40.
Checking: 40 years old is 25 years older than 15 years old.
In 10 years, the father will be 50 and Rana will be 25 and 50 is
twice 25. So the ansewr is correct.
Edwin